OpenStudy (anonymous):

Can anyone help me with this problem?: R(r1+r2)=r1r2 solve the formula for r2

OpenStudy (anonymous):

Rr1 + Rr2 = r1r2 r1r2 - Rr2 = Rr1 r2 = Rr1/(r1-r2)

OpenStudy (anonymous):

can you tell me how you got to steps 2 and 3 please?